# Tri-resonant leptogenesis from modular symmetry neutrino models

### v1

@@ 4 Study for modular ${\rm S}_{4}$ symmetry model
One can see that this model also contains 8 real parameters in total: $\alpha v_{d}$, $\beta/\alpha$, $\gamma/\alpha$, $g^{2}v^{2}_{u}/M_{0}$, $|g^{\prime}/g|$, ${\rm arg}(g^{\prime}/g)$, ${\rm Re}(\tau)$ and ${\rm Im}(\tau)$. The numerical calculation shows that this model can be consistent with the neutrino oscillation experimental results. For $M_{\rm D}$ in Eq. (34), the model can be consistent with the neutrino oscillation experimental results only in the NO case, and the numerical search finds two pairs of distinct local minima of $\Delta\chi^{2}$ corresponding to two pairs of distinct values of $\tau$. The two minima in each pair lead to opposite values of the Dirac and Majorana phases, but the same values of all other observables. The cases belonging to the first pair are denoted as Case A, while those belonging to the second pair are denoted as Case B. For these two cases, the resulting predictions for the best-fit values and the allowed ranges of the input parameters and observables are collected in Table 3 [22]. For $M_{\rm D}$ in Eq. (35), the model can be consistent with the neutrino oscillation experimental results in both the NO and IO cases. In the IO case, the numerical search finds two pairs of distinct local minima of $\Delta\chi^{2}$ corresponding to two pairs of distinct values of $\tau$, and the cases belonging to these two pairs are respectively denoted as Case C and Case D. In the NO case, the numerical search only finds one pair of local minima of $\Delta\chi^{2}$ corresponding to one pair of distinct values of $\tau$, and the cases belonging to this pair are denoted as Case E. For these three cases, the resulting predictions for the best-fit values and the allowed ranges of the input parameters and observables are collected in Table 4 [22].
@@
${\rm Im}(\tau)$ 1.01 $1.004-1.018$ 1.005 $0.9988-1.008$
$\beta/\alpha$ 9.465 $7.693-12.39$ 0.03306 $0.02529-0.04074$
$\gamma/\alpha$ 0.002205 $0.001941-0.002472$ 0.0001307 $0.0000982-0.0001663$
${\rm Re}(g^{\prime}/g)$ 0.233 $-0.02544-0.4417$ 0.4097 $0.3241-0.5989$
@@
$m_{e}/m_{\mu}$ 0.004802 $0.00422-0.005383$ 0.004802 $0.004211-0.005384$
$m_{\mu}/m_{\tau}$ 0.0565 $0.04317-0.06961$ 0.05649 $0.04318-0.06962$
$\sin^{2}\theta_{12}$ 0.305 $0.2656-0.3449$ 0.305 $0.2662-0.3455$
$\sin^{2}\theta_{13}$ 0.02125 $0.01912-0.02383$ 0.02125 $0.01914-0.02383$
$\sin^{2}\theta_{23}$ 0.551 $0.4838-0.5999$ 0.551 $0.4322-0.601$
$m_{1}$ [eV] 0.01746 $0.01185-0.02143$ 0.02074 $0.01918-0.02428$
$m_{2}$ [eV] 0.01945 $0.01473-0.02307$ 0.02244 $0.02101-0.02574$
$m_{3}$ [eV] 0.05288 $0.05075-0.05452$ 0.05406 $0.05314-0.05577$
@@
Table 3: For the model given in section 4 with $M_{\rm D}$ in Eq. (34), in the phenomenologically viable NO case, the predictions for the best-fit values and the allowed ranges of the input parameters and observables in Case A and Case B. These results are taken from Tables 5a-5b in Ref. [22].
@@
${\rm Im}(\tau)$ 1.523 $1.088-1.594$ 1.397 $1.236-1.529$ 1.309 $1.246-1.385$
$\beta/\alpha$ 17.82 $9.32-23.66$ 15.35 $10.79-21.09$ 0.000243 $0.0002004-0.0002864$
$\gamma/\alpha$ 0.003243 $0.00227-0.003733$ 0.002924 $0.002443-0.003459$ 0.03335 $0.02799-0.03926$
${\rm Re}(g^{\prime}/g)$ $-0.8714$ $-(0.7956-1.148)$ $-1.32$ $-(1.131-1.447)$ $-0.06454$ $-(0.01697-0.1215)$
@@
$m_{e}/m_{\mu}$ 0.004797 $0.004215-0.005378$ 0.004786 $0.004221-0.005386$ 0.004797 $0.004393-0.005197$
$m_{\mu}/m_{\tau}$ 0.05655 $0.04348-0.0698$ 0.0554 $0.04343-0.06968$ 0.05626 $0.04741-0.0654$
$\sin^{2}\theta_{12}$ 0.303 $0.2657-0.3436$ 0.3031 $0.2657-0.3436$ 0.311 $0.2895-0.3375$
$\sin^{2}\theta_{13}$ 0.02175 $0.01957-0.0242$ 0.02184 $0.01954-0.0242$ 0.02185 $0.02041-0.02351$
$\sin^{2}\theta_{23}$ 0.5571 $0.4551-0.6026$ 0.5577 $0.5482-0.6013$ 0.4469 $0.43-0.4614$
$m_{1}$ [eV] 0.0513 $0.04882-0.05207$ 0.05122 $0.05023-0.05212$ 0.01774 $0.01703-0.01837$
$m_{2}$ [eV] 0.05201 $0.04958-0.05274$ 0.05193 $0.05098-0.05279$ 0.0197 $0.01906-0.02025$
$m_{3}$ [eV] 0.01512 $0.00316-0.0163$ 0.01495 $0.01223-0.01649$ 0.05299 $0.05251-0.05346$
@@
Table 4: For the model given in section 4 with $M_{\rm D}$ in Eq. (35), in their respective phenomenologically viable neutrino mass ordering cases, the predictions for the best-fit values and the allowed ranges of the input parameters and observables in Case C, Case D and Case E. These results are taken from Tables 5c-5e in Ref. [22].
Now, in a similar way as in section 3, we study the consequences of these five cases for tri-resonant leptogenesis in the scenarios that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10) and generated from the RGE effects, respectively. For the former scenario, the left panel of Figure 8 has shown the total CP asymmetries $\varepsilon_{\rm tot}$ in the decays of $N_{I}$ as functions of $2\mu/\Gamma_{\rm tot}$. We see that the results in these five cases are very similar to one another, and the CP asymmetries can take their maximally allowed values around $2\mu/\Gamma_{\rm tot}\sim 1$. This will be helpful for us to reproduce the observed value of $\eta_{\rm B}$. The right panel of Figure 8 has shown the values of $\mu$ that accommodate a successful leptogenesis as functions of $M_{0}$. The results for the five cases are very similar to one another, and the ratios of $\mu$ to $M_{0}$ are around $10^{-9}$ or $10^{-18}$ at $M_{0}=10^{3}$ GeV. For the latter scenario, the left panel of Figure 9 has shown the total CP asymmetries $\varepsilon_{\rm tot}$ in the decays of $N_{I}$ as functions of $M_{0}$, while the right panel of Figure 9 has shown the generated $\eta_{\rm B}$ as functions of $M_{0}$. These results have also been obtained by taking $\Lambda_{\rm FS}=10^{16}$ GeV. The results show that only for Case E and $M_{0}\sim 10^{12}$ GeV can the observed value of $\eta_{\rm B}$ be successfully reproduced, while the the magnitudes of $\eta_{\rm B}$ are always larger than its observed value in the other four cases. It is interesting to recall that Case E holds in the NO case which is slightly preferred by the current experimental results compared to the IO case [2, 3].
Figure 8: For the modular ${\rm S}_{4}$ symmetry model given in section 4, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10), the total CP asymmetries $\varepsilon_{\rm tot}$ in the decays of $N_{I}$ as functions of $2\mu/\Gamma_{\rm tot}$ (left panel), and the values of $\mu$ that accommodate a successful leptogenesis as functions of $M_{0}$ (right panel).
Figure 9: For the modular ${\rm S}_{4}$ symmetry model given in section 4, in the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects, the total CP asymmetries $\varepsilon_{\rm tot}$ in the decays of $N_{I}$ as functions of $M_{0}$ (left panel), and the generated $\eta_{\rm B}$ as functions of $M_{0}$ (right panel).
@@ 5 Study for modular ${\rm A}_{5}$ symmetry model
$M_{\rm D}=y_{0}v_{u}\left(\begin{array}[]{ccc}\sqrt{3}Y_{1}&Y_{4}&Y_{3}\\ Y_{5}&-\sqrt{2}Y_{3}&-\sqrt{2}Y_{2}\\ Y_{2}&-\sqrt{2}Y_{5}&-\sqrt{2}Y_{4}\end{array}\right)\;.$ (40)
@@
Similar to the models in sections 3 and 4, in the present model the charged-lepton masses are essentially controlled by $\alpha v_{d}$, $\beta/\alpha$ and $\gamma/\alpha$, while the neutrino masses and mixing angles are mainly governed by $y^{2}_{0}v^{2}_{u}/M_{0}$, $\tau$, $\varphi_{2}$ and $\varphi_{3}$. The number of parameters is further reduced by enforcing the CP conservation which restricts all the Yukawa couplings together with $\varphi_{2}$ and $\varphi_{3}$ to be real (with the VEV of $\tau$ serving as the only source of CP violation). The numerical calculation shows that this model can be consistent with the neutrino oscillation experimental only in the NO case and it further predicts $m_{1}=0$. The resulting predictions for the best-fit values of the input parameters and observables are collected in Table 5.
Re($\tau$) Im$(\tau)$ $\varphi_{2}$ $\varphi_{3}$ $\alpha\cos\beta\cdot 10^{3}$ $\beta/\alpha$ $\gamma/\alpha$ $y^{2}_{0}\sin^{2}\beta$/$M_{0}$ (eV-1) --0.3615 0.2412 0.04759 0.3731 3.368 1.310 --0.2413 0.0001639
$m_{1}\cdot 10^{2}~{}{\rm eV}^{-1}$ $m_{2}\cdot 10^{2}~{}{\rm eV}^{-1}$ $m_{3}\cdot 10^{2}~{}{\rm eV}^{-1}$ $\sin^{2}\theta_{12}$ $\sin^{2}\theta_{13}$ $\sin^{2}\theta_{23}$ $\delta/\pi$ $(\alpha_{21}$-$\alpha_{31})/\pi$ $\chi^{2}_{\textrm{min}}$ 0 0.860 5.02 0.292 0.0228 0.449 1.63 1.68 11.1
@@
Table 5: For the model given in section 5, in the phenomenologically viable NO case, the predictions for the best-fit values of the input parameters and observables. These results are taken from Tables 8 in Ref. [24].
For this model, in a similar way as in sections 3 and 4, we study the consequences of it for tri-resonant leptogenesis in the scenarios that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10) and generated from the RGE effects, respectively. For the former scenario, the left panel of Figure 10 has shown the values of $\mu$ that accommodate a successful leptogenesis as functions of $M_{0}$. The results show that the ratio of $\mu$ to $M_{0}$ is around $10^{-8}$ or $10^{-10}$ at $M_{0}=10^{3}$ GeV. For the latter scenario, the right panel of Figure 10 has shown the generated $\eta_{\rm B}$ as functions of $M_{0}$. These results have also been obtained by taking $\Lambda_{\rm FS}=10^{16}$ GeV. The results show that the observed value of $\eta_{\rm B}$ can not be successfully reproduced.
Figure 10: For the modular ${\rm A}_{5}$ symmetry model given in section 5, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10), the values of $\mu$ that accommodate a successful leptogenesis as functions of $M_{0}$ (left panel). In the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects, the generated $\eta_{\rm B}$ as functions of $M_{0}$ (right panel).

### v3

@@ 5 Study for modular ${\rm S}_{4}$ symmetry model
One can see that this model also contains 8 real parameters in total: $\alpha v_{d}$, $\beta/\alpha$, $\gamma/\alpha$, $g^{2}v^{2}_{u}/M_{0}$, $|g^{\prime}/g|$, ${\rm arg}(g^{\prime}/g)$, ${\rm Re}(\tau)$ and ${\rm Im}(\tau)$. The numerical calculation shows that this model can be consistent with the neutrino oscillation experimental results. For $M_{\rm D}$ in Eq. (50), the model can be consistent with the neutrino oscillation experimental results only in the NO case, and the numerical search finds two pairs of distinct local minima of $\Delta\chi^{2}$ corresponding to two pairs of distinct values of $\tau$. The two minima in each pair lead to opposite values of the Dirac and Majorana phases, but the same values of all other observables. The cases belonging to the first pair are denoted as Case A, while those belonging to the second pair are denoted as Case B. For these two cases, the resulting predictions for the best-fit values and the allowed ranges of the input parameters and observables are collected in Table 2 here [29]. For $M_{\rm D}$ in Eq. (51), the model can be consistent with the neutrino oscillation experimental results in both the NO and IO cases. In the IO case, the numerical search finds two pairs of distinct local minima of $\Delta\chi^{2}$ corresponding to two pairs of distinct values of $\tau$, and the cases belonging to these two pairs are respectively denoted as Case C and Case D. In the NO case, the numerical search only finds one pair of local minima of $\Delta\chi^{2}$ corresponding to one pair of distinct values of $\tau$, and the cases belonging to this pair are denoted as Case E. For these three cases, the resulting predictions for the best-fit values and the allowed ranges of the input parameters and observables are collected in Table 3 here [29].
@@
${\rm Im}(\tau)$ 1.01 1.0041.0181.004-1.018 1.005 0.99881.0080.9988-1.008
$\beta/\alpha$ 9.465 7.69312.397.693-12.39 0.03306 0.025290.040740.02529-0.04074
$\gamma/\alpha$ 0.002205 0.0019410.0024720.001941-0.002472 0.0001307 0.00009820.00016630.0000982-0.0001663
${\rm Re}(g^{\prime}/g)$ 0.233 0.025440.4417-0.02544-0.4417 0.4097 0.32410.59890.3241-0.5989
@@
$m_{e}/m_{\mu}$ 0.004802 0.004220.0053830.00422-0.005383 0.004802 0.0042110.0053840.004211-0.005384
$m_{\mu}/m_{\tau}$ 0.0565 0.043170.069610.04317-0.06961 0.05649 0.043180.069620.04318-0.06962
$\sin^{2}\theta_{12}$ 0.305 0.26560.34490.2656-0.3449 0.305 0.26620.34550.2662-0.3455
$\sin^{2}\theta_{13}$ 0.02125 0.019120.023830.01912-0.02383 0.02125 0.019140.023830.01914-0.02383
$\sin^{2}\theta_{23}$ 0.551 0.48380.59990.4838-0.5999 0.551 0.43220.6010.4322-0.601
$m_{1}$ [eV] 0.01746 0.011850.021430.01185-0.02143 0.02074 0.019180.024280.01918-0.02428
$m_{2}$ [eV] 0.01945 0.014730.023070.01473-0.02307 0.02244 0.021010.025740.02101-0.02574
$m_{3}$ [eV] 0.05288 0.050750.054520.05075-0.05452 0.05406 0.053140.055770.05314-0.05577
$\sum_{i}m_{i}$ [eV] 0.0898 $[0.07735,0.09887]$ 0.09724 $[0.0935,0.1056]$
$m_{ee}$ [eV] 0.01699 $[0.01177,0.02002]$ 0.01983 $[0.01847,0.02275]$
@@
Table 2: For the modular ${\rm S}_{4}$ symmetry model in section 5 with $M_{\rm D}$ in Eq. (50), in the phenomenologically viable NO case, the predictions for the best-fit values and the allowed ranges of the input parameters and observables in Case A and Case B. These results are taken from Tables 5a-5b in Ref. [29].
@@
${\rm Im}(\tau)$ 1.523 1.0881.5941.088-1.594 1.397 1.2361.5291.236-1.529 1.309 1.2461.3851.246-1.385
$\beta/\alpha$ 17.82 9.3223.669.32-23.66 15.35 10.7921.0910.79-21.09 0.000243 0.00020040.00028640.0002004-0.0002864
$\gamma/\alpha$ 0.003243 0.002270.0037330.00227-0.003733 0.002924 0.0024430.0034590.002443-0.003459 0.03335 0.027990.039260.02799-0.03926
${\rm Re}(g^{\prime}/g)$ 0.8714-0.8714 (0.79561.148)-(0.7956-1.148) 1.32-1.32 (1.1311.447)-(1.131-1.447) 0.06454-0.06454 (0.016970.1215)-(0.01697-0.1215)
@@
$m_{e}/m_{\mu}$ 0.004797 0.0042150.0053780.004215-0.005378 0.004786 0.0042210.0053860.004221-0.005386 0.004797 0.0043930.0051970.004393-0.005197
$m_{\mu}/m_{\tau}$ 0.05655 0.043480.06980.04348-0.0698 0.0554 0.043430.069680.04343-0.06968 0.05626 0.047410.06540.04741-0.0654
$\sin^{2}\theta_{12}$ 0.303 0.26570.34360.2657-0.3436 0.3031 0.26570.34360.2657-0.3436 0.311 0.28950.33750.2895-0.3375
$\sin^{2}\theta_{13}$ 0.02175 0.019570.02420.01957-0.0242 0.02184 0.019540.02420.01954-0.0242 0.02185 0.020410.023510.02041-0.02351
$\sin^{2}\theta_{23}$ 0.5571 0.45510.60260.4551-0.6026 0.5577 0.54820.60130.5482-0.6013 0.4469 0.430.46140.43-0.4614
$m_{1}$ [eV] 0.0513 0.048820.052070.04882-0.05207 0.05122 0.050230.052120.05023-0.05212 0.01774 0.017030.018370.01703-0.01837
$m_{2}$ [eV] 0.05201 0.049580.052740.04958-0.05274 0.05193 0.050980.052790.05098-0.05279 0.0197 0.019060.020250.01906-0.02025
$m_{3}$ [eV] 0.01512 0.003160.01630.00316-0.0163 0.01495 0.012230.016490.01223-0.01649 0.05299 0.052510.053460.05251-0.05346
$\sum_{i}m_{i}$ [eV] 0.1184 0.1020.12080.102-0.1208 0.1181 0.11390.12120.1139-0.1212 0.09043 0.088740.091950.08874-0.09195
$m_{ee}$ [eV] 0.0263 0.022880.045510.02288-0.04551 0.0310 0.025150.036770.02515-0.03677 0.006967 0.0064820.0072880.006482-0.007288
@@
Table 3: For the modular ${\rm S}_{4}$ symmetry model in section 5 with $M_{\rm D}$ in Eq. (51), in their respective phenomenologically viable neutrino mass ordering cases, the predictions for the best-fit values and the allowed ranges of the input parameters and observables in Case C, Case D and Case E. These results are taken from Tables 5c-5e in Ref. [29].
Figure 8: For the five cases of modular ${\rm S}_{4}$ symmetry model in section 5, the total CP asymmetries $\varepsilon_{T}$ for the decays of $N_{I}$ as functions of $M_{0}$ for the benchmark values of $\tan\beta=10$ (left panel) and 3030 (right panel).
Figure 9: For the five cases of modular ${\rm S}_{4}$ symmetry model in section 5, the resulting $\eta_{\rm B}$ as functions of $M_{0}$ for the benchmark values of $\tan\beta=10$ (left panel) and 3030 (right panel). The horizontal gray dashed line shows the observed value of $\eta_{\rm B}$.
Figure 10: For the five cases of modular ${\rm S}_{4}$ symmetry model in section 5, the values of $M_{0}$ versus $\tan\beta$ that allow for a successful reproduction of the observed value of $\eta_{\rm B}$.
Now, we study the RGE induced tri-resonant leptogenesis for these five cases. Figure 8 has shown the total CP asymmetries $\varepsilon_{T}$ for the decays of $N_{I}$ as functions of $M_{0}$ for the benchmark values of $\tan\beta=10$ and 30. For $M_{0}<10^{12}$ GeV, all of the CP asymmetries are almost independent of $M_{0}$, while they change very sharply for $M_{0}>10^{12}$ GeV. Figure 9 has shown the resulting $\eta_{\rm B}$ as functions of $M_{0}$ also for the benchmark values of $\tan\beta=10$ and 30. On can see that the observed value of $\eta_{\rm B}$ has chance to be reproduced successfully for appropriate values of $M_{0}$ and $\tan\beta$. This can be seen more clearly from Figure 10 which has shown the values of $M_{0}$ versus $\tan\beta$ that allow for a successful reproduction of the observed value of $\eta_{\rm B}$. The results show that for Case A and Case C there may exist two allowed values of $M_{0}$ for each allowed value of $\tan\beta$. The results show that for Case A and Case C there is an upper bound on $\tan\beta$ (about 45 and 28, respectively) and for each allowed value of $\tan\beta$ there may exist two allowed values of $M_{0}$. Importantly, Case C allows relatively low values of $M_{0}$, which may be reachable for future direct measurements. For the other cases, the allowed values of $M_{0}$ (about $10^{14}$, $10^{9}$ and $10^{13}$ GeV for Case B, Case D and Case E, respectively) are almost independent of $\tan\beta$.
@@ 6 Study for modular ${\rm A}_{5}$ symmetry model
$M_{\rm D}=y_{0}v_{u}\left(\begin{array}[]{ccc}\sqrt{3}Y_{1}&Y_{5}&Y_{2}\\ Y_{4}&-\sqrt{2}Y_{3}&-\sqrt{2}Y_{5}\\ Y_{3}&-\sqrt{2}Y_{2}&-\sqrt{2}Y_{4}\end{array}\right)\;.$ (56)
@@
Similar to the models in sections 4 and 5, in the present model the charged-lepton masses are essentially controlled by $\alpha v_{d}$, $\beta/\alpha$ and $\gamma/\alpha$, while the neutrino masses and mixing angles are mainly governed by $y^{2}_{0}v^{2}_{u}/M_{0}$, $\tau$, $\varphi_{2}$ and $\varphi_{3}$. The number of parameters is further reduced by enforcing the CP conservation which restricts all the Yukawa couplings together with $\varphi_{2}$ and $\varphi_{3}$ to be real (with the VEV of $\tau$ serving as the only source of CP violation). The numerical calculation shows that this model can be consistent with the neutrino oscillation experimental results only in the NO case and it further predicts $m_{1}=0$. The resulting predictions for the best-fit values of the input parameters and observables are collected in Table 4 here.
[…]
$m_{1}({\rm meV})$ $m_{2}({\rm meV})$ $m_{3}({\rm meV})$ $m_{ee}({\rm meV})$ $\sin^{2}\theta_{12}$ $\sin^{2}\theta_{13}$ $\sin^{2}\theta_{23}$ $\delta/\pi$ $(\alpha_{21}$-$\alpha_{31})/\pi$ $\chi^{2}_{\textrm{min}}$ 0 8.60 50.2 1.3 0.292 0.0228 0.449 1.63 1.68 11.1
@@
Table 4: For the modular ${\rm A}_{5}$ symmetry model in section 6, in the phenomenologically viable NO case, the predictions for the best-fit values of the input parameters and observables. These results are taken from Table 8 in Ref. [31].
Figure 11: For the modular ${\rm A}_{5}$ symmetry model in section 6, the resulting CP asymmetries for the decays of $N_{I}$ as functions of $M_{0}$ for the benchmark values of $\tan\beta=$10 (left panel) and 30 (right panel).
Figure 12: For the modular ${\rm A}_{5}$ symmetry model in section 6, the resulting $\eta_{\rm B}$ as functions of $M_{0}$ for the benchmark values of $\tan\beta=$10, 20, 30 and 50 (left panel). The horizontal gray dashed line shows the observed value of $\eta_{\rm B}$. the values of $M_{0}$ versus $\tan\beta$ that allow for a successful reproduction of the observed value of $\eta_{\rm B}$ (right panel).
Now, we study the RGE induced tri-resonant leptogenesis for the above model. Figure 11 has shown the resulting CP asymmetries for the decays of $N_{I}$ as functions of $M_{0}$ for the benchmark values of $\tan\beta=10$ and 30. For $M_{0}<10^{12}$ GeV, all of the CP asymmetries are almost independent of $M_{0}$, while they change very sharply for $M_{0}>10^{12}$ GeV. The left panel of Figure 12 has shown the resulting $\eta_{\rm B}$ as functions of $M_{0}$ for some benchmark values of $\tan\beta$. The result shows that the resulting $\eta_{\rm B}$ is almost independent of $M_{0}$ when $M_{0}$ takes low values, and varies rapidly when $M_{0}$ becomes sufficiently large. Finally, the right panel of Figure 12 has shown the values of $M_{0}$ versus $\tan\beta$ that allow for a successful reproduction of the observed value of $\eta_{\rm B}$. The results show that the allowed values of $M_{0}$ can vary between $10^{7}$ and $10^{10}$ GeV for $\tan\beta$ from 10 to 50.
Before proceeding to the summary section, we briefly discuss the potential constraints imposed by charged lepton violating processes on the models considered in this work. Among such processes, the most stringent limit stems from the $\mu\to e\gamma$ decay, as reported in Ref. [33]:
${\mathrm{B}r}\left(\mu\to e\gamma\right)<3.1\times 10^{-13}\left(90\%\;{\mathrm{C.L.}}\right)\;.$ (58)
Within the type-I seesaw framework, the branching ratio for $\mu\to e\gamma$ is given by [34]:
${\mathrm{B}r}\left(\mu\to e\gamma\right)=\frac{2\alpha_{\rm em}}{32\pi}\left|\sum^{3}_{i=1}\left(RV\right)^{*}_{\mu i}\left(RV\right)_{ei}\left[G\left(X\right)-G\left(0\right)\right]\right|^{2}\;,$ (59)
where $X\equiv(M_{0}/M_{W})^{2}$ with $M_{1,2,3}\simeq M_{0}$, and $G\left(X\right)$ is a loop integration function ranging in value from 4/3 to 10/3. Here, $V$ denotes the unitary matrix that diagonalizes the Majorana mass matrix of heavy right-handed neutrinos, and $R$ is determined by $R^{*}\simeq M_{\rm D}M^{-1}$. This branching ratio can be approximated as:
${\mathrm{B}r}\left(\mu\to e\gamma\right)\simeq\frac{2\alpha_{\rm em}}{32\pi}\left(\frac{m_{\nu}}{M_{0}}\right)^{2}\;.$ (60)
For the heavy neutrino mass range considered in this paper ($10^{3}$ to $10^{15}$ GeV), the branching ratio lies on the order of $10^{-51}$ to $10^{-27}$, which is far below the $10^{-13}$ upper limit from the latest MEG II experiment.